OpenStudy (cheesecakekitten):

WILL MEDAL AND FAN!!! Which shows solutions to x + 4y = 9 if x and y must be whole numbers? A. [(1, 2), (5, 1), (9, 0)] B. [(0, 9), (1, 4), (2, 1)] C. [(2, 1), (1, 6), (0, 9)] D. [(0, 9), (1, 5), (2, 1)]

OpenStudy (smilie2015):

A

OpenStudy (smilie2015):

If u want reason i can explain

OpenStudy (smilie2015):

(1,2) x+4y=9 1+(4*2)=9 1+8=9 9=9 (5,1) 5+(4*1)=9 9=9 (9,0) 9+(4.0)=9 9+0=9 9=0

OpenStudy (cheesecakekitten):

THANKS A THOUSAND

OpenStudy (cheesecakekitten):

What is the slope of a line perpendicular to the line with equation y = 5x - 2? A. 5 B. -1/5 C. -2 D. -5

OpenStudy (smilie2015):

B